Wang Co. manufactures and sells a single product that sells for $300 per unit; variable costs are $174. Annual fixed costs are $852,600. Current sales volume is $4,230,000. Compute the break-even point in units.

Wang Co. manufactures and sells a single product that sells for $250 per unit; variable costs are $145 per unit. Annual fixed costs are $873,600. Current sales volume is $4,280,000. Management targets an annual pre-tax income of $1,205,000. Compute the unit sales to earn the target pre-tax net income.

Wang Co. manufactures and sells a single product that sells for $640 per unit; variable costs are $352 per unit. Annual fixed costs are $985,500. Current sales volume is $4,390,000. Compute the current margin of safety in dollars
Answer #1
Dear Student
Thank you for referring Chegg
Please give thumbs up if you are satisfied
Statement showing Computations
Particulars Amount
Sales price per unit                      300.00
Less Variable Expenses per unit                   (174.00)
Contribution Margin per unit                      126.00
Fixed expenses             852,600.00
break-even point in units= 852,600/126                        6,767
b)
Sales price per unit                      250.00
Less Variable Expenses per unit                   (145.00)
Contribution Margin per unit                      105.00
Fixed expenses             873,600.00
annual pre-tax income          1,205,000.00
Contribution desired          2,078,600.00
unit sales to earn the target pre-t+A4975ax net income. = 2078600/105                      19,796
c)
Sales price per unit                      640.00
Less Variable Expenses per unit                   (352.00)
Contribution Margin per unit                      288.00
CM Ratio = 288/640 45.00%
Fixed expenses             985,500.00
break-even point in $ =985500/45%                2,190,000
Current sales volume          4,390,000.00
current margin of safety in dollars = 4390,000 - 2190,000          2,200,000.00